# Graph Max Shift: A Hill-Climbing Method for Graph Clustering

### v1

@@ 6 Proofs
As we noted earlier on the in paper, the arguments supporting Theorem 3.3 are similar to those establishing the consistency of Max Shift in [2], but not exactly the same as the flat kernel (3.2) in use here is not twice differentiable or 2nd order, and the arguments in [2] rely crucially on these properties being satisfied by the kernel function used to produce a density estimate. The reason for that is that the proof structure is laid out in [2] to be applicable not just to Max Shift but to other methods such as Mean Shift. Because we focus on (Graph) Max Shift, we are able to tailor the arguments to the particular situation we are confronted with — and actually simplify them in the process.
@@ 6.1.1 Density
$\kappa_{0}=\sup_{x\in\mathbb{R}^{d}}f(x),\qquad\kappa_{1}=\sup_{x\in\mathbb{R}^{d}}\|\nabla f(x)\|,\qquad\kappa_{2}=\sup_{x\in\mathbb{R}^{d}}\|\nabla^{2}f(x)\|,$ (6.1)
@@ 6.1.2 Upper level sets
Let $x^{*}$ be a mode of $f$ and let $s^{*}=f(x^{*})$. There exists $\lambda_{1}\geq\lambda_{0}>0$ and $\delta>0$ such that, for any $x\in\bar{B}(x^{*},\delta)$, the eigenvalues of $\nabla^{2}f(x)$ belong to $[-\lambda_{1},-\lambda_{0}]$. Also, for any $s\in(s^{*}-\frac{1}{2}\lambda_{0}\delta^{2},s^{*})$,
@@
The first part of the statement relies on the fact that $\nabla^{2}f(x^{*})$ is negative definitive and that $\nabla^{2}f$ is continuous. The second part is a restatement of Lemma 5.6 in [4]. ∎
@@ 6.1.3 Approximations
For the stochastic term on the right-hand side of (6.9), from the main result in [13] and how $\epsilon=\epsilon_{n}$ is assumed to vary, we have that
@@
which quantifies how dense the sample $\mathcal{Y}$ is in the set $\mathcal{U}_{s}$. The following is standard. (See [2, Lem 5.6] for a detailed proof that follows the standard arguments and applies almost unchanged to the present context.)
@@
We recall that the graph distance between two nodes is the minimum number of hops, or length of the shortest path, to travel between the two nodes. Let $H_{ij}$ denote the graph distance between nodes $i$ and $j$. Under some circumstances, in the context of a geometric graph, the graph distance tracks the Euclidean distance. The following relies on a result established in [1].
@@
For any $s>0$, there is a constant $C>0$ such that, if $\alpha_{s/2}/\epsilon\leq 1/C$,
$0\leq H_{ij}\epsilon-\|y_{i}-y_{j}\|\leq C(\alpha_{s/2}/\epsilon)\|y_{i}-y_{j}\|+\epsilon,$
@@
Let $\tilde{H}_{ij}$ denote the graph distance between nodes $i$ and $j$ within the subgraph defined by the nodes $k$ such that $y_{k}\in\mathcal{S}$. In such circumstances, if $\alpha_{s/2}/\epsilon\leq 1/4$, [1, Th 1] yields
@@
6.2 Auxiliary results
The first few results are on the behavior of a typical hill-climbing sequence built by Graph Max Shift, namely, when started at $y_{0}\in\mathcal{Y}$,
$y_{k+1}\in\operatorname*{arg\,max}\limits_{y\in\mathcal{Y}\cap\bar{B}(y_{k},\epsilon)}\hat{f}_{\epsilon}(y),\quad k\geq 0.$ (6.15)
@@ 6.2 Auxiliary results
There is $C_{0}>0$ such that, provided $\kappa_{1}\epsilon+2\eta\leq s/2$ and $\alpha_{s/2}\leq\epsilon$, for the Graph Max Shift sequence $(y_{k})$ initialized at any $y_{0}\in\mathcal{U}_{s}$,
$\epsilon-C_{0}\frac{(\epsilon^{2}+\alpha_{s/2}+\eta)}{\|\nabla f(y_{k})\|}\leq\|y_{k+1}-y_{k}\|\leq\epsilon,\quad\forall k\geq 0.$ (6.16)
@@
The upper bound is by construction of the sequence. The lower bound follows from applying the Cauchy–Schwarz inequality to the inner product in the left-hand side of (6.28) below. (Note that only the upper bound in (6.16) is used in the establishing of that inequality.) ∎
Lemma 6.7.
Under the same setting as Lemma 6.6,
$y_{k+1}-y_{k}=\|y_{k+1}-y_{k}\|N(y_{k})\pm 2C_{0}\frac{\epsilon^{1/2}(\epsilon^{2}+\alpha_{s/2}+\eta)^{1/2}}{\|\nabla f(y_{k})\|^{1/2}},\quad\forall k\geq 0.$ (6.17)
Proof.
Let $\alpha$ be short for $\alpha_{s/2}$. Let $y_{k+1}^{\prime}$ be the closest point in $\mathcal{Y}$ to $x_{k+1}^{\prime}:=y_{k}+(\epsilon-\alpha)N(y_{k})$ and assume $\alpha\leq\epsilon$, so that $\|x_{k+1}^{\prime}-y_{k}\|=\epsilon-\alpha\leq\epsilon$. We have
$f(x_{k+1}^{\prime})$ $\geq f(y_{k})-\kappa_{1}\epsilon$ (6.18)
$\geq\hat{f}(y_{k})-\kappa_{1}\epsilon-\eta$ (6.19)
$\geq\hat{f}(y_{0})-\kappa_{1}\epsilon-\eta$ (6.20)
$\geq f(y_{0})-\kappa_{1}\epsilon-2\eta$ (6.21)
$\geq s-\kappa_{1}\epsilon-2\eta,$ (6.22)
using (6.2), (6.8), the fact that $\hat{f}(y_{k})$ is increasing by definition, and (6.8) again. Therefore, provided $\kappa_{1}\epsilon+2\eta\leq s/2$, we have $f(x_{k+1}^{\prime})\geq s/2$, so that $x_{k+1}^{\prime}\in\mathcal{U}_{s/2}$, and by definition of $\alpha$, $\|y_{k+1}^{\prime}-x_{k+1}^{\prime}\|\leq\alpha$. Therefore,
$\|y_{k+1}^{\prime}-y_{k+1}\|\leq\|y_{k+1}^{\prime}-x_{k+1}^{\prime}\|+\|x_{k+1}^{\prime}-y_{k}\|\leq\epsilon-\alpha+\alpha\leq\epsilon,$
by the triangle inequality, so that, by construction of $y_{k+1}$,
$\hat{f}(y_{k+1})\geq\hat{f}(y_{k+1}^{\prime}).$
Continuing, using (6.8), (6.2), and (6.4), we derive
$\hat{f}(y_{k+1}^{\prime})$ $\geq f(y_{k+1}^{\prime})-\eta$ (6.23)
$\geq f(x_{k+1}^{\prime})-\kappa_{1}\alpha-\eta$ (6.24)
$\geq f(y_{k})+\nabla f(y_{k})^{\top}(x_{k+1}^{\prime}-y_{k})-\kappa_{2}\|x_{k+1}^{\prime}-y_{k}\|^{2}-\kappa_{1}\alpha-\eta$ (6.25)
$=f(y_{k})+\|\nabla f(y_{k})\|(\epsilon-\alpha)-\kappa_{2}(\epsilon-\alpha)^{2}-\kappa_{1}\alpha-\eta$ (6.26)
$\geq f(y_{k})+\|\nabla f(y_{k})\|\epsilon-\kappa_{2}\epsilon^{2}-2\kappa_{1}\alpha-\eta.$ (6.27)
Similarly, in the other direction,
$\hat{f}(y_{k+1})\leq f(y_{k+1})+\eta\leq f(y_{k})+\nabla f(y_{k})^{\top}(y_{k+1}-y_{k})+\kappa_{2}\epsilon^{2}+\eta.$
Combining these gives
$f(y_{k})+\|\nabla f(y_{k})\|\epsilon-\kappa_{2}\epsilon^{2}-2\kappa_{1}\alpha-\eta\leq f(y_{k})+\nabla f(y_{k})^{\top}(y_{k+1}-y_{k})+\kappa_{2}\epsilon^{2}+\eta,$
@@
$N(y_{k})^{\top}(y_{k+1}-y_{k})\geq\epsilon-\frac{2\big{(}\kappa_{2}\epsilon^{2}+\kappa_{1}\alpha+\eta\big{)}}{\|\nabla f(y_{k})\|}.$ (6.28)
Let $u_{k}:=y_{k+1}-y_{k}$ and $\xi_{k}:=2(\kappa_{2}\epsilon^{2}+\kappa_{1}\alpha+\eta)/\|\nabla f(y_{k})\|$. If $\epsilon-\xi_{k}<0$, we simply write
$\big{\|}u_{k}-\|u_{k}\|N(y_{k})\big{\|}\leq 2\|u_{k}\|\leq 2\epsilon\leq 2(\epsilon\xi_{k})^{1/2}.$ (6.29)
If $\epsilon-\xi_{k}\geq 0$, we may write
$\big{\|}u_{k}-\|u_{k}\|N(y_{k})\big{\|}^{2}$ $=2\|u_{k}\|^{2}-2\|u_{k}\|N(y_{k})^{\top}u_{k}$ (6.30)
@@
Either way, (6.17) holds. ∎
@@
For any $s>0$, there is $\delta>0$ and $C_{1}>0$ such that, if $\epsilon,\eta\leq 1/C_{1}$, for any mode $x^{*}\in\mathcal{U}_{s}$, the Graph Max Shift sequence initialized at a point $y_{0}\in\mathcal{Y}\cap B(x^{*},\delta)$ converges to a point in $\mathcal{Y}$ within distance $C_{1}(\epsilon+(\alpha_{s/2}+\eta)/\epsilon)$ of $x^{*}$.
@@
Because there are finitely many modes in $\mathcal{U}_{s}$, it is enough to consider a single mode $x^{*}\in\mathcal{U}_{s}$. For that mode, let $s^{*}:=f(x^{*})$, so that $s^{*}\geq s$, and let $\delta_{1},\lambda_{0},\lambda_{1}$ be as defined in Lemma 6.1. Then, by a Taylor development of $\nabla f$,
$\|\nabla f(x)\|\geq\lambda_{0}\|x-x^{*}\|,\quad\text{for any $x\in B(x^{*},\delta_{1})$.}$ (6.34)

### v2

@@ 6 Proofs
As we noted earlier on the in paper, the arguments supporting Theorem 3.3 are similar to those establishing the consistency of Max Shift in Arias-Castro and Qiao (2025), but not exactly the same as the flat kernel (3.2) in use here is not twice differentiable or 2nd order, and the arguments in Arias-Castro and Qiao (2025) rely crucially on these properties being satisfied by the kernel function used to produce a density estimate. The reason for that is that the proof structure is laid out in Arias-Castro and Qiao (2025) to be applicable not just to Max Shift but to other methods such as Mean Shift. Because we focus on (Graph) Max Shift, we are able to tailor the arguments to the particular situation we are confronted with — and actually simplify them in the process.
@@ 6.1.1 Density
$\kappa_{0}=\sup_{x\in\mathbb{R}^{d}}f(x),\hskip 18.49988pt\kappa_{1}=\sup_{x\in\mathbb{R}^{d}}\|\nabla f(x)\|,\hskip 18.49988pt\kappa_{2}=\sup_{x\in\mathbb{R}^{d}}\|\nabla^{2}f(x)\|,$ (6.1)
@@ 6.1.2 Upper level sets
Let $x^{*}$ be a mode of $f$ and let $s^{*}=f(x^{*})$. There exists $\lambda_{1}\geq\lambda_{0}>0$ and $\delta>0$ such that, for any $x\in\bar{B}(x^{*},\delta)$, the eigenvalues of $\nabla^{2}f(x)$ belong to $\lambda_{1},-\lambda_{0}]$. Also, for any $s\in(s^{*}-\frac{1}{2}\lambda_{0}\delta^{2},s^{*})$,
@@
The first part of the statement relies on the fact that $\nabla^{2}f(x^{*})$ is negative definitive and that $\nabla^{2}f$ is continuous. The second part is a restatement of Lemma 5.6 in Arias-Castro and Qiao (2023a). ∎
@@ 6.1.3 Approximations
For the stochastic term on the right-hand side of (6.9), from (Giné and Guillou, 2002, Th 2.1) and because $\epsilon\gg(\log(n)/n)^{1/d}$ in Theorem 3.3, we have
@@
which quantifies how dense the sample $\mathcal{Y}$ is in the set $\mathcal{U}_{s}$. The following is standard. (See (Arias-Castro and Qiao, 2025, Lem 5.6) for a detailed proof that follows the standard arguments and applies almost unchanged to the present context.)
@@
We recall that the graph distance between two nodes is the minimum number of hops, or length of the shortest path, to travel between the two nodes. Let $H_{ij}$ denote the graph distance between nodes $i$ and $j$. Under some circumstances, in the context of a geometric graph, the graph distance tracks the Euclidean distance. The following relies on a result established in Arias-Castro et al. (2021).
@@
For any $s>0$, there is a constant $A_{0}>0$ such that, if $\alpha_{s/2}/\epsilon\leq 1/A_{0}$,
$0\leq H_{ij}\epsilon-\|y_{i}-y_{j}\|\leq A_{0}(\alpha_{s/2}/\epsilon)\|y_{i}-y_{j}\|+\epsilon,$
@@
Let $\tilde{H}_{ij}$ denote the graph distance between nodes $i$ and $j$ within the subgraph defined by the nodes $k$ such that $y_{k}\in\mathcal{S}$. In such circumstances, if $\alpha_{s/2}/\epsilon\leq 1/4$, (Arias-Castro et al., 2021, Th 1) yields
@@
6.2 Results about Max Shift
$x_{k+1}\in\operatorname*{arg\,max}\big\{f(x):x\in\bar{B}(x_{k},\epsilon)\big\},\qquad k\geq 0.$ (6.15)
The first one is essentially Lem 3.3 in Arias-Castro and Qiao (2025), but not quite, while the second one is a straightforward consequence of the first.
@@ 6.2 Results about Max Shift
For the Max Shift sequence $(x_{k})$ initialized at any $x_{0}$,
$x_{k+1}-x_{k}=\|x_{k+1}-x_{k}\|N(x_{k})\pm\frac{2(2\kappa_{2})^{1/2}\epsilon^{3/2}}{\|\nabla f(x_{k})\|^{1/2}}.$ (6.16)
@@
Define $\tilde{x}_{k+1}:=x_{k}+\epsilon N(x_{k})$. Using (6.4), we derive
$f(\tilde{x}_{k+1})$ $\geq f(x_{k})+\nabla f(x_{k})^{\top}(\tilde{x}_{k+1}-x_{k})-\kappa_{2}\|\tilde{x}_{k+1}-x_{k}\|^{2}$ (6.17)
$=f(x_{k})+\|\nabla f(x_{k})\|\epsilon-\kappa_{2}\epsilon^{2}.$ (6.18)
Similarly, going in the other direction with starting point the fact that $\tilde{x}_{k+1}\in\bar{B}(x_{k},\epsilon)$,
$f(\tilde{x}_{k+1})\leq f(x_{k+1})\leq f(x_{k})+\nabla f(x_{k})^{\top}(x_{k+1}-x_{k})+\kappa_{2}\epsilon^{2}.$
Combining these inequalities, we obtain
$f(x_{k})+\|\nabla f(x_{k})\|\epsilon-\kappa_{2}\epsilon^{2}\leq f(x_{k})+\nabla f(x_{k})^{\top}(x_{k+1}-x_{k})+\kappa_{2}\epsilon^{2},$
@@
$N(x_{k})^{\top}(x_{k+1}-x_{k})\geq\epsilon-\frac{2\kappa_{2}\epsilon^{2}}{\|\nabla f(x_{k})\|}.$ (6.19)
Let $u_{k}:=x_{k+1}-x_{k}$ and $\xi_{k}:=2\kappa_{2}\epsilon^{2}/\|\nabla f(x_{k})\|$, so that (6.19) reads $N(x_{k})^{\top}u_{k}\geq\epsilon-\xi_{k}.$ If $\epsilon<\xi_{k}$, we have
$\big\|u_{k}-\|u_{k}\|N(x_{k})\big\|\leq 2\|u_{k}\|\leq 2\epsilon\leq 2(\epsilon\xi_{k})^{1/2};$ (6.20)
while if $\epsilon\geq\xi_{k}$, we have
$\big\|u_{k}-\|u_{k}\|N(x_{k})\big\|^{2}$ $=2\|u_{k}\|^{2}-2\|u_{k}\|N(x_{k})^{\top}u_{k}$ (6.21)
@@
Either way,
$u_{k}-\|u_{k}\|N(x_{k})=\pm 2(\epsilon\xi_{k})^{1/2}=\frac{2(2\kappa_{2})^{1/2}\epsilon^{3/2}}{\|\nabla f(x_{k})\|^{1/2}},$ (6.25)
so that (6.16) holds. ∎
@@
For the Max Shift sequence $(x_{k})$ initialized at any $x_{0}$,
$\|x_{k+1}-x_{k}\|\geq\epsilon-\frac{2\kappa_{2}\epsilon^{2}}{\|\nabla f(x_{k})\|}.$ (6.26)
@@
The bound follows from applying the Cauchy–Schwarz inequality to the inner product on the left-hand side of (6.19). ∎
Note that combining (6.16) and (6.26), we have
$x_{k+1}-x_{k}$ $=\epsilon N(x_{k})\pm\left(\frac{2(2\kappa_{2})^{1/2}\epsilon^{3/2}}{\|\nabla f(x_{k})\|^{1/2}}+\frac{2\kappa_{2}\epsilon^{2}}{\|\nabla f(x_{k})\|}\right)$ (6.27)
$=\epsilon N(x_{k})\pm\frac{2(2\kappa_{2})^{1/2}\kappa_{1}^{1/2}\epsilon^{3/2}+2\kappa_{2}\epsilon^{2}}{\|\nabla f(x_{k})\|},$ (6.28)
using the fact that $\|\nabla f(x_{k})\|\leq\kappa_{1}$.
The third result is distilled from the proof of Th 3.1–3.2 and arguments provided in Sec 3.2.1 in the same paper Arias-Castro and Qiao (2025).
Lemma 6.8.
For any $p>0$, there is $s>0$ and a subset $\Omega\subset\mathcal{U}_{s}$ with $\int_{\Omega}f\geq 1-p$ such that any point in $\Omega$ is in the basin of attraction of a mode. Moreover, for any $\delta>0$, there is $A_{1}>0$ such that, if $\epsilon\leq 1/A_{1}$, starting at any point $x_{0}\in\Omega$ in the basin of attraction of some mode $x^{*}$, the sequence $(x_{k})$ that Max Shift computes enters $B(x^{*},\delta)$ in at most $A_{1}/\epsilon$ steps, while $\|\nabla f(x_{k})\|\geq 1/A_{1}$ for all $k\geq 0$ such that $x_{k}\notin B(x^{*},\delta)$.
With (6.28), in the context of Lemma 6.8,
$x_{k+1}-x_{k}$ $=\epsilon N(x_{k})\pm C_{0}\epsilon^{3/2},$ (6.29)
for all $k\geq 0$ such that $x_{k}\notin B(x^{*},\delta)$, where $C_{0}:=2(2\kappa_{2})^{1/2}\kappa_{1}^{1/2}A_{1}+2\kappa_{2}A_{1}^{1/2}$.
6.3 Proof of Theorem 3.3
We start by proving that, with arbitrarily high probability, a point which is in the basin of attraction of a mode is moved to a terminal point near the mode. In fact, we show that this happens in a uniform manner within the density upper level sets. From that, the proof of Theorem 3.3 will easily follow.
Proposition 6.9.
For any $p>0$, there is $s>0$ and a subset $\Omega\subset\mathcal{U}_{s}$ with $\int_{\Omega}f\geq 1-p$, and a constant $A$ such that any data point in $\Omega$ in the basin of attraction of a mode is moved by Graph Max Shift to a point within distance $A(\epsilon+(\alpha_{s/2}+\eta)/\epsilon)$ of the mode.
Proof of Proposition 6.9.
Fix $p>0$, let $s>0$ and $\Omega\subset\mathcal{U}_{s}$ be as in Lemma 6.8. Let $\alpha$ be short for $\alpha_{s/2}$. Because $\mathcal{U}_{s}$ is compact, it is enough to prove the result for $\epsilon$ and $q:=(\alpha+\eta)/\epsilon$ small enough.
Take any data point $y^{0}\in\mathcal{Y}\cap\Omega$ as a starting point and let $(y^{k})$ denote the sequence in (3.4) computed by Graph Max Shift; let $(x_{k})$ denote the sequence computed by Max Shift based on $f$ as defined in (6.15) with the same starting point $x_{0}=y^{0}$. The following lemmas are proved later.
Lemma 6.10.
There is $A_{2}>0$ such that, provided $\kappa_{1}\epsilon+2\eta\leq s/2$ and $\alpha_{s/2}\leq\epsilon$, for the Graph Max Shift sequence $(y^{k})$ initialized at any $y^{0}\in\mathcal{U}_{s}$,
$y^{k+1}-y^{k}=\|y^{k+1}-y^{k}\|N(y^{k})\pm A_{2}\frac{\epsilon^{1/2}(\epsilon^{2}+\alpha_{s/2}+\eta)^{1/2}}{\|\nabla f(y^{k})\|^{1/2}},$ (6.30)
and
$\|y^{k+1}-y^{k}\|\geq\epsilon-A_{2}\frac{(\epsilon^{2}+\alpha_{s/2}+\eta)}{\|\nabla f(y^{k})\|}.$ (6.31)
When $\epsilon$ and $q$ are small enough, the conditions of Lemma 6.10 are met. In that case, combining (6.30) and (6.31), we have
$y^{k+1}-y^{k}$ $=\epsilon N(y^{k})\pm\left(A_{2}\frac{\epsilon^{2}+\alpha+\eta}{\|\nabla f(y^{k})\|}+A_{2}\frac{\epsilon^{1/2}(\epsilon^{2}+\alpha+\eta)^{1/2}}{\|\nabla f(y^{k})\|^{1/2}}\right)$ (6.32)
$=\epsilon N(y^{k})\pm C_{1}\frac{q^{1/2}\epsilon}{\|\nabla f(y^{k})\|},$ (6.33)
for some constant $C_{1}$, using the fact that $\|\nabla f(y^{k})\|\leq\kappa_{1}$.
Lemma 6.11.
For any $s>0$, there is $\delta>0$ and $A_{3}>0$ such that, for any mode $x^{*}\in\mathcal{U}_{s}$, the Graph Max Shift sequence initialized at a point $y^{0}\in\mathcal{Y}\cap B(x^{*},\delta)$ converges to a point in $\mathcal{Y}$ within distance $A_{3}(\epsilon+(\alpha_{s/2}+\eta)/\epsilon)$ of $x^{*}$.
When $\epsilon$ and $q$ are small enough, the conditions of Lemma 6.11 are met. In that case, let $x^{*}$ be the mode to which $y^{0}$ is attracted by gradient ascent and let $\delta$ be as in Lemma 6.11. By Lemma 6.11, it suffices to show that
$y^{k}\in\bar{B}(x^{*},\delta)$ for some $k$. (6.34)
Assuming that $x_{0}=y^{0}\notin B(x^{*},\delta/2)$, for otherwise our job is done, let $k_{\#}$ be such that $x_{k_{\#}}\in B(x^{*},\delta/2)$ while $x_{k}\notin B(x^{*},\delta/2)$ for $k<k_{\#}$, which is well-defined by Lemma 6.8. The same lemma tells us that $k_{\#}\leq A_{1}/\epsilon$. Assume that $q$ is small enough that $q\leq\delta/2$, so that $\alpha\leq\delta/2$. Then, by the triangle inequality, to establish (6.34) it is enough to show that
$\|x_{k_{\#}}-y^{k_{\#}}\|\leq\delta/2.$ (6.35)
Define $C_{2}=2A_{1}C_{1}$ and $C_{3}=(2A_{1}+1)\kappa_{2}$, and assume $\epsilon$ and $q$ are small enough that
$(C_{0}\epsilon^{1/2}+C_{2}q^{1/2})C_{3}^{-1}(\exp(C_{3}A_{1})-1)\leq\frac{1}{2A_{1}\kappa_{2}}\wedge\frac{\delta}{2}.$ (6.36)
We show by recursion that
$\|x_{k}-y^{k}\|\leq(C_{0}\epsilon^{1/2}+C_{2}q^{1/2})C_{3}^{-1}(\exp(C_{3}\epsilon k)-1),$ (6.37)
for all $k\leq k_{\#}$. We start the recursion by noting that (6.37) is true at $k=0$ since $x_{0}=y^{0}$. Now, assume (6.37) holds at some $0\leq k<k_{\#}$. By the fact that $k_{\#}\leq A_{1}/\epsilon$ and (6.36), we also have
$\|x_{k}-y^{k}\|\leq\frac{1}{2A_{1}\kappa_{2}}.$ (6.38)
Then, by Lemma 6.8 and a Taylor development,
$\|\nabla f(y^{k})\|\geq\|\nabla f(x_{k})\|-\kappa_{2}\|x_{k}-y^{k}\|\geq\frac{1}{A_{1}}-\kappa_{2}\frac{1}{2A_{1}\kappa_{2}}=\frac{1}{2A_{1}}.$
Combining this with (6.33), we find that
$y^{k+1}-y^{k}$ $=\epsilon N(y^{k})\pm C_{2}q^{1/2}\epsilon.$ (6.39)
The inequalities (6.29) and (6.39), together with the triangle inequality, imply
$\|x_{k+1}-y^{k+1}\|$ $\leq\|x_{k}-y^{k}\|+\epsilon\|N(x_{k})-N(y^{k})\|+C_{0}\epsilon^{3/2}+C_{2}q^{1/2}\epsilon.$ (6.40)
In view of (6.6), and the fact (justified by the same Taylor development) that $\|\nabla f(z)\|\geq 1/2A_{1}$ for all $z$ on the line segment joining $x_{k}$ and $y^{k}$,
$\|N(x_{k})-N(y^{k})\|\leq C_{3}\|x_{k}-y^{k}\|,$
implying that
$\|x_{k+1}-y^{k+1}\|$ $\leq(1+C_{3}\epsilon)\|x_{k}-y^{k}\|+C_{0}\epsilon^{3/2}+C_{2}q^{1/2}\epsilon.$ (6.41)
Using the recursion hypothesis (6.37), we find after some simplifications based on $e^{a}-1-a\geq 0$ for all $a$, that
$\|x_{k+1}-y^{k+1}\|\leq(C_{0}\epsilon^{1/2}+C_{2}q^{1/2})C_{3}^{-1}(\exp(C_{3}\epsilon(k+1))-1),$ (6.42)
so that the recursion carries on. (The arguments above underly a discrete form of Grönwall’s inequality.) In particular, we have established that
$\|x_{k_{\#}}-y^{k_{\#}}\|\leq(C_{0}\epsilon^{1/2}+C_{2}q^{1/2})C_{3}^{-1}(\exp(C_{3}A_{1})-1),$ (6.43)
which in view of (6.36) implies (6.35) . ∎
Proof of Lemma 6.10.
Let $\alpha$ be short for $\alpha_{s/2}$. Let $\tilde{y}^{k+1}$ be the closest point in $\mathcal{Y}$ to $\tilde{x}_{k+1}:=y^{k}+(\epsilon-\alpha)N(y^{k})$ and assume $\alpha\leq\epsilon$, so that $\|\tilde{x}_{k+1}-y^{k}\|=\epsilon-\alpha\leq\epsilon$. We have
$f(\tilde{x}_{k+1})$ $\geq f(y^{k})-\kappa_{1}\epsilon$ (6.44)
$\geq\hat{f}(y^{k})-\kappa_{1}\epsilon-\eta$ (6.45)
$\geq\hat{f}(y^{0})-\kappa_{1}\epsilon-\eta$ (6.46)
$\geq f(y^{0})-\kappa_{1}\epsilon-2\eta$ (6.47)
$\geq s-\kappa_{1}\epsilon-2\eta,$ (6.48)
using (6.2), (6.8), the fact that $\hat{f}(y^{k})$ is increasing by definition, and (6.8) again. Therefore, provided $\kappa_{1}\epsilon+2\eta\leq s/2$, we have $f(\tilde{x}_{k+1})\geq s/2$, so that $\tilde{x}_{k+1}\in\mathcal{U}_{s/2}$, and by definition of $\alpha$, $\|\tilde{y}^{k+1}-\tilde{x}_{k+1}\|\leq\alpha$. Therefore,
$\|\tilde{y}^{k+1}-y^{k+1}\|\leq\|\tilde{y}^{k+1}-\tilde{x}_{k+1}\|+\|\tilde{x}_{k+1}-y^{k+1}\|\leq\epsilon-\alpha+\alpha\leq\epsilon,$
by the triangle inequality, so that, by construction of $y^{k+1}$,
$\hat{f}(y^{k+1})\geq\hat{f}(\tilde{y}^{k+1}),$
and continuing, using (6.8), (6.2), and (6.4), we derive
$\hat{f}(y^{k+1})\geq\hat{f}(\tilde{y}^{k+1})$ $\geq f(\tilde{y}^{k+1})-\eta$ (6.49)
$\geq f(\tilde{x}_{k+1})-\kappa_{1}\alpha-\eta$ (6.50)
$\geq f(y^{k})+\nabla f(y^{k})^{\top}(\tilde{x}_{k+1}-y^{k})-\kappa_{2}\|\tilde{x}_{k+1}-y^{k}\|^{2}-\kappa_{1}\alpha-\eta$ (6.51)
$=f(y^{k})+\|\nabla f(y^{k})\|(\epsilon-\alpha)-\kappa_{2}(\epsilon-\alpha)^{2}-\kappa_{1}\alpha-\eta$ (6.52)
$\geq f(y^{k})+\|\nabla f(y^{k})\|\epsilon-\kappa_{2}\epsilon^{2}-2\kappa_{1}\alpha-\eta.$ (6.53)
In the other direction, using (6.8) and (6.4),
$\hat{f}(y^{k+1})\leq f(y^{k+1})+\eta\leq f(y^{k})+\nabla f(y^{k})^{\top}(y^{k+1}-y^{k})+\kappa_{2}\epsilon^{2}+\eta.$
Combining these inequalities gives
$f(y^{k})+\|\nabla f(y^{k})\|\epsilon-\kappa_{2}\epsilon^{2}-2\kappa_{1}\alpha-\eta\leq f(y^{k})+\nabla f(y^{k})^{\top}(y^{k+1}-y^{k})+\kappa_{2}\epsilon^{2}+\eta,$
which after some simplifications yields
$N(y^{k})^{\top}(y^{k+1}-y^{k})\geq\epsilon-\frac{2\big(\kappa_{2}\epsilon^{2}+\kappa_{1}\alpha+\eta\big)}{\|\nabla f(y^{k})\|}.$ (6.54)
We then obtain (6.30) from (6.54) exactly as we obtained (6.16) from (6.19) in the proof of Lemma 6.6, but this time with $u_{k}:=y^{k+1}-y^{k}$ and $\xi_{k}:=2(\kappa_{2}\epsilon^{2}+\kappa_{1}\alpha+\eta)/\|\nabla f(y^{k})\|$.
We then derive (6.31) by applying the Cauchy–Schwarz inequality to the inner product on the left-hand side of (6.54). ∎
Proof of Lemma 6.11.
@@
Because $\mathcal{U}_{s}$ is compact, it is enough to prove the result when $\epsilon$ and $(\alpha_{s/2}+\eta)/\epsilon$ are small enough. And since there are finitely many modes in $\mathcal{U}_{s}$, it is enough to consider a single mode $x^{*}\in\mathcal{U}_{s}$. For that mode, let $s^{*}:=f(x^{*})$, so that $s^{*}\geq s$, and let $\delta_{1},\lambda_{0},\lambda_{1}$ be as defined in Lemma 6.1. Then, by a Taylor development of $\nabla f$,
$\|\nabla f(x)\|\geq\lambda_{0}\|x-x^{*}\|,\qquad\text{for any $x\in B(x^{*},\delta_{1})$.}$ (6.55)